What is the acceleration of the car?

I thought it would be -9.8m/s^2 since it's free fall, but the answer is -8.94 m/s^2.
Answer #1

The car is not in free fall. Free fall refers to the condition when there is only earth's gravity acting on a falling object and no other forces are considered.

Here you have to consider the air resistance.

The problem states that the car goes to 60 mph from 0 mph in 3s, which means it goes from 0 m/s to 26.8224 m/s in 3 s, ---(60 mph = 26.8224 m/s)

using , v = u + at

-26.8224 m/s = 0 + a(3s) --- (we have taken downward to be negative)

or, a = -8.94 m/s2

***************************************************************************************************
Check the answer and let me know immediately if you find something wrong... I will rectify the mistakes asap if any